Which is a hyperbola? Thank you SO much! D:
4x² - 4y = 16 or 4x + 4y² = 16

Found 2 solutions by rapaljer, richwmiller:
Answer by rapaljer(4671)

Neither of these is a hyperbola! A hyperbola is usually in the form like
x^2 - y^2 = 16 or y^2 - x^2 = 16.

ADDITIONAL NOTE: In order to be a hyperbola in this form, it must have BOTH an x^2 and a y^2 with opposite sign for coefficients. There are exceptions to this when you get to rotated hyperbolas (this is probably beyond the scope of your question), for example, xy = 1 or y=1/x. However, I don't think this is relevant to this question. As the other tutor mentioned, these are both parabolas, since you have an x^2 OR a y^2, but NOT both. I hope this helps. If you have further questions, feel free to send me an Email at rapaljer@seminolestate.edu.

One more note: If your teacher says that one of these is a hyperbola, then you must have copied the problem wrong!! For example, if the question had been 4x^2-4y^2=16, then THAT would be a hyperbola!

Dr. Robert J. Rapalje, Retired

Answer by richwmiller(17219)

They are both parabolas.